OpenStudy (anonymous):

2/7x=10. To get x by itself , what do i do?

OpenStudy (callisto):

Divide both sides by 2/7 to isolate x.

OpenStudy (anonymous):

Firstly multiply by 7 both the sides...

OpenStudy (anonymous):

@ela099 are you listening what we said ??

OpenStudy (anonymous):

uh yes

OpenStudy (anonymous):

thanks

hero (hero):

\(\frac{2x}{7} = \frac{10}{1}\) Cross multiply to get: 2x = 70 Divide both sides by 2 to get \( x = \frac{70}{2}\) x = 35